OpenStudy (anonymous):

Can someone help? (9/100)^2/^3 Please show your work

OpenStudy (anonymous):

Please help!

OpenStudy (phi):

is this \[ \left( \frac{9}{100}\right)^\frac{2}{3} \]?

OpenStudy (phi):

or maybe \[ \left( \frac{9}{100}\right)^\frac{3}{2} \] ??

OpenStudy (phi):

or something else ?